OpenStudy (anonymous):

find sec(-7π/3) without a calculator

OpenStudy (idku):

can you find the cosine of -7pi/3 ?

OpenStudy (idku):

-7pi/3 = -420 = -360 - 60 then use the cosine (a+b) rule

OpenStudy (anonymous):

\[\sec \left( \frac{ -7 \pi }{ 3 } \right)=\sec \left( \frac{ 7 \pi }{ 3 } \right)=\sec \left( 2 \pi+\frac{ \pi }{ 3 } \right)\] \[=\sec \frac{ \pi }{ 3 }=\frac{ 1 }{ \cos \frac{ \pi }{ 3 } }=\frac{ 1 }{ \frac{ 1 }{ 2 } }=?\]